OpenStudy (anonymous):

How many permutations exist of the letters a, b, c, d taken two at a time? 12 8 2

ganeshie8 (ganeshie8):

taking two at a time ? you can take them like this right ? ab, ba ac, ca ad, da bc, cb bd, db cd, dc total how many are they ?

OpenStudy (anonymous):

6?

Parth (parthkohli):

Easier way to do this is to calculate 4C2.

Parth (parthkohli):

\( \color{Black}{\Rightarrow {4! \over (4 - 2)!} = {4 \times 3 \times 2 \over 2} }\)

terenzreignz (terenzreignz):

Are you sure it isn't 4P2? It was permutations after all...

OpenStudy (anonymous):

i mean 12

ganeshie8 (ganeshie8):

it is 4p2. im trying to make him see what the question is asking for...

Parth (parthkohli):

oops 4P2

terenzreignz (terenzreignz):

@ParthKohli I think what you gave was 4P2 after all 4C2 looks like \[\frac{4!}{(4-2)!2!}\]

Parth (parthkohli):

I said that wrong haha

Parth (parthkohli):

Yeah I misspelt x{

OpenStudy (anonymous):

im sooooooooo confused lol

terenzreignz (terenzreignz):

That one crucial letter, eh? :)

Parth (parthkohli):

something else was in my mind hehe

Parth (parthkohli):

I answered a combinations question a while ago

terenzreignz (terenzreignz):

hey @sdk, you want it to be explained thoroughly? I see that I'm drifting away from the main question, lol sorry